B will take 3 days longer than A to complete a piece of work. If they together can complete the work in two days. Find the time each alone will take to complete the work.

Answer by richwmiller(17219)


b=a+3
(2/a)+(2/(a+3))=1
a=3
b=6
check
2/3+2/6=1
2/3+1/3=1
1=1
ok